when 3 multiplied by 5 less than a number, the result is 9 less than 5 times the number. Find the numbers
Answer by Cromlix(4381)

3 multiplied by 5 less than a number, the result is 9 less than 5 times the number.
3(x - 5) = 5x - 9
3x - 15 = 5x - 9
3x - 5x = -9 + 15
-2x = 6
x = -3
The number is -3
Proof:
3(x - 5) = 5x - 9
3(-3 - 5) = 5(-3) - 9
3(-8) = -5 - 9
-24 = -24
Hope this helps.
:-)